What is the x-coordinate of this equation? : y=x-7 y=-2x+5
Answer by Alan3354(69443)

What is the x-coordinate of this equation? : y=x-7 y=-2x+5
------------
If you mean of the 2 equations:
y=x-7
y=-2x+5
Since they both equal y:
x-7 = -2x+5
3x = 12
x = 4
---
y = -3